OpenStudy (anonymous):

My answer is A What is the solution to the equation 9 -3x ≈ 7 ? (1 points) Answer x = 0.376 x = 0.295 x = -0.295 x = -0.376

OpenStudy (campbell_st):

so is this question \[9^{-3x} = 7\]

OpenStudy (anonymous):

yes sir

OpenStudy (campbell_st):

ok... so the log law you keep needing is \[logx^a = a \times \log x\] so using this and taking the log of both sides \[\log(9^{-3x}) = \log(7)\] using the above log law \[-3x \times \log(9) = \log(7)\] divide both sides of the equation by log(9) \[-3x = \frac{\log(7)}{\log(9)}\] now divide both sides of the equation by -3 \[x = \frac{-1}{3} \times \frac{\log(7)}{\log(9)}\] if you put that into a calculator you should get one of your answer choices.

OpenStudy (anonymous):

are you serious? that's really all you have to do?

OpenStudy (campbell_st):

to make it simple input \[-1 \div 3 \times \log(7) \div \log(9)\]

OpenStudy (anonymous):

I'm going to look at your work so i stop asking questions about this

OpenStudy (campbell_st):

yep... thats the log law you seem to have been using in the last few questions.

OpenStudy (anonymous):

how did you know to put the -3x on the outside of the equation?

OpenStudy (anonymous):

Also, the answer is C :)

OpenStudy (campbell_st):

because its the power.....there are a few log laws you need to know.. these notes are quite good... if you scroll down you'll see the log laws http://tutorial.math.lamar.edu/Classes/CalcI/LogFcns.aspx

OpenStudy (anonymous):

thank you for the material

OpenStudy (campbell_st):

the answer is C

OpenStudy (anonymous):

youre a life saver! thank you so very much, really

OpenStudy (campbell_st):

glad to help... logs are a pain